OpenStudy (anonymous):

Simplify: 4! 24 6 120 240

jimthompson5910 (jim_thompson5910):

A similar problem: 5! = 5*4*3*2*1 = 120 So using this idea, what is 4!

OpenStudy (anonymous):

4! = 4(3)(2)(1) = 12*2 = 24

OpenStudy (anonymous):

factorial (!) means you take the number*the number - 1*the number - 2 and so forth until you reach 3*2*1

OpenStudy (anonymous):

so 12! would be 12*11*10*9*8*7*6*5*4*3*2*1

OpenStudy (anonymous):

Simplify: 7/3 210 24 840 6

OpenStudy (anonymous):

do you mean 7! / 3! because that's 7(6)(5)(4) = 840.

OpenStudy (anonymous):

YEA THATS WHAT I MEAN

OpenStudy (anonymous):

ok. well, it's 840.